OpenStudy (anonymous):

HELP PLEASE! ( -6ab^3 / 5a ) * ( 10b/-24ab^2 )

OpenStudy (anonymous):

\[\frac{ -6ab^3 }{ 5a } \times \frac{ 10b }{ -24ab^2 }\]

OpenStudy (anonymous):